OpenStudy (anonymous):

1/2(x - 3) + (3/2 - x) = 5x

OpenStudy (anonymous):

\[\frac{1}{2(x-3)} + \frac{3}{2-x} =5\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

it has to be this one because the one written is incorrect and the only solution is for x=0

OpenStudy (anonymous):

no - it is 1/2 * (x - 3) + (3/2 - x) = 5x

OpenStudy (anonymous):

then x=0

OpenStudy (anonymous):

multiply everything by 2

OpenStudy (anonymous):

x-3+3-x=10x

OpenStudy (anonymous):

x=0 as stated before

OpenStudy (anonymous):

why multiply everything by 2? to cancel out the fraction yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you both!

OpenStudy (anonymous):

you're welcome